Consider four electric charges at the corners of a
Answer #1

F1 = K*Q2*Q1 / a^2 = (9*10^9*10*10^-9*20*10^-9) / (2*10^-6)^2
=450000 N

F2 = K*Q1*Q3 / 2a^2 = (9*10^9*6*10^-9*10*10^-9) / (2*(2*10^-6)^2)
= 67500 N

F3 = K*Q1*Q4 / a^2 = (9*10^9*10*10^-9*-8*10^-9) / (2*10^-6)^2
= 180000 N

F = (F2Cos45 + F3)x + (-F1 + F2Sin45)y
= (67500cos45 + 180000)x + (-450000 + 67500sin45)y
= 227729.70Nx - 402270.29Ny

Magitude F = sqrt(Fx^2 + Fy^2)

= 462257.7 N